OpenStudy (anonymous):

Solve 2x2 - 12x + 20 = 0 I know I'll have to factor it, and that it will end with an imaginary number as well, any help???

OpenStudy (jdoe0001):

\(\large 2x^2 - 12x + 20 = 0\implies \begin{array}{llll} 2(x^2&-6x&+10)=0\\ &\quad \uparrow &\quad \uparrow \\ &-3-2&-3\cdot -2 \end{array}\)

OpenStudy (jdoe0001):

hmmm wait.... shoot.. ahemm -3-2 is -5... ok

OpenStudy (jdoe0001):

hehhe

OpenStudy (anonymous):

well, here are the options : 3 ± i 3 ± 2i 1 ± 2i 2 ± 3i

OpenStudy (anonymous):

which one would it be ?

OpenStudy (jdoe0001):

have you covered the quadratic formula yet?

OpenStudy (anonymous):

yeah, do I need to use that to solve it then ? (if so, how ? )

OpenStudy (jdoe0001):

\(\bf 2x^2 - 12x + 20 = 0\implies2(x^2-6x+10)=0\implies x^2-6x+10=0 \\ \quad \\ \textit{quadratic formula}\\ {\color{blue}{ 1}}x^2{\color{red}{ -6}}x{\color{green}{ +10}} \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)